OpenStudy (anonymous):

what is the radius of a circle with the area of 36in^2

TheSmartOne (thesmartone):

Hi there! Welcome to OpenStudy

TheSmartOne (thesmartone):

Do you know the formula for area of a circle?

OpenStudy (alexandervonhumboldt2):

sue formula area=r^2*pi

OpenStudy (alexandervonhumboldt2):

*use

OpenStudy (alexandervonhumboldt2):

32in^2=pi*r^2 pi=3.14

OpenStudy (alexandervonhumboldt2):

substitute value of pi and solve for r

OpenStudy (alexandervonhumboldt2):

wopps

OpenStudy (alexandervonhumboldt2):

i meant 36in^2=r^2*pi

OpenStudy (anonymous):

so it would equal to 3.39

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

you are correct

TheSmartOne (thesmartone):

@Infiniti03 The formula for finding the area of a circle is \(\sf Area = \pi r^2\) Now, we are given the area as \(\sf 36\) \(\sf\Large 36 = \pi\cdot r^2\) So we need to sovle for r.

OpenStudy (anonymous):

thanks

OpenStudy (alexandervonhumboldt2):

Since you are new here read the tutorial of how to use openstudy: http://openstudy.com/study#/updates/5655f2d1e4b0af360a7cfaa3

OpenStudy (anonymous):

thanks